A 25Ω resistor is wired directly across a 120-Vac source. What happens to the time-average power, rms voltage, and rms current when the resistor’s value is changed to 50 Ω?
Answer #1

When we change resistance by 50Ohm then Vrms will remain same but Irms and Pavg will change as follows

Vrms = Vo/Sqrt(2) = 120/1.414 = 84.86 V

Irms = Vrms / R = 84.86/ 50 = 1.697 A

Pavg = Vrms x Irms = 144.007 Watt